OpenStudy (anonymous):

Solve open parentheses 8 x plus 1 close parentheses to the 1 half power equals 5

OpenStudy (anonymous):

(8x+1)^1/2=5 this can be written as (8x+1) whole square root take square on both sides to eliminate 1/2 ( square root) it will be (8x+1)2/2=5^2 8x+1=25 8x=24 x=3 :D